OpenStudy (anonymous):

Oregon is about 400 miles from west to east, and 300 miles from north to south. If a map of Oregon is 15 inches tall (from north to south), about how wide is the map?

jimthompson5910 (jim_thompson5910):

solve for x 15/300 = x/400

jimthompson5910 (jim_thompson5910):

let me know what you get

OpenStudy (wolf1728):

Or to put it another way North - South / East - West 300 / 400 = 15 inches / x